# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ 4.2 Two-matrix weighted bounds for paraproducts
We now prove that $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ Using Lemma 5.3 of [DKPS2024] we have
@@
$b_{R}^{\varepsilon}\coloneq|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|,\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E},$
and
$\widetilde{M}_{\bm{\mathcal{D}},U}f\coloneq\sup_{R\in\bm{\mathcal{D}}}\langle|\mathcal{U}_{R}f|\rangle_{R}\mathbf{1}_{R}.$
By the well-known unweighted bounds for paraproducts in the scalar setting (see e. g. [Blasco_Pott_2005]) we have
@@
By definition, $\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}=\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ From [DKPS2024, Proposition 4.1] we also have
$\|\widetilde{M}_{\bm{\mathcal{D}},U}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim 1.$
It follows that
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}$ $\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\widetilde{M}_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}$
$\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|f\|_{\mathrm{L}^{p}(U)}.$
@@
$\|\Pi_{\bm{\mathcal{D}},B}^{(00)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}$ $=\|\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}\|_{L^{p^{\prime}}(V^{\prime})\rightarrow L^{p^{\prime}}(U^{\prime})}\sim\|B^{\ast}\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(V^{\prime},U^{\prime},p^{\prime})}$
$\sim_{p,d}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
We adapt the factorization trick from the proof of [Holmes_Petermichl_Wick_2018, Proposition 6.1]. We have
$(\Gamma_{\bm{\mathcal{D}},B}f,g)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\langle B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\varepsilon},g_{R}^{\delta}\rangle.$
@@
$|(\Gamma_{\bm{\mathcal{D}},B}f,g)|$ $=\bigg{|}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}\text{tr}((B_{R}^{\varepsilon})^{\ast}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T})\bigg{|}$
@@
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R|}}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T},\quad R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \varepsilon\in\mathcal{E}.$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{L^{p}(U)}\|g\|_{L^{p^{\prime}}(V^{\prime})}.$
We have
$\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\bigg{|}V(x)^{-1/p}\bigg{(}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\bigg{)}\mathcal{U}_{R}\bigg{|}^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}|^{2}\cdot|\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$=2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\leavevmode\nobreak\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq 2^{n+m}\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}.$
It is proved in [DKPS2024, Lemma 5.2 and Corollary 5.4] that
$\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{L^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},\qquad\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{L^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{L^{p^{\prime}}(V^{\prime})},$
@@
Next, following the terminology of Holmes–Petermichl–Wick [Holmes_Petermichl_Wick_2018, Subsection 6.1] we define the following so-called “mixed” biparameter paraproducts acting on (suitable) $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}$:
$\Pi_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(h_{R_{1}}^{\varepsilon_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}},\quad\Pi_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\varepsilon_{2}}\right)B_{R}^{\varepsilon}\langle f^{\varepsilon_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}h_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{1\oplus\varepsilon_{2}\oplus\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\\ \\ delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}h_{R}^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{1},1}_{R_{1}}\rangle_{R_{2}},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{1\oplus\varepsilon_{1}\oplus\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{\delta_{1},\varepsilon_{2}}_{R}.$
It is clear that $(\Pi_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(10)}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense. Observe also that a change of summation variables yields
$\Gamma_{\bm{\mathcal{D}},B}^{(10),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}\left(\frac{\mathbf{1}_{R_{1}}}{|R_{1}|}\otimes h_{R_{2}}^{\delta_{2}}\right)B_{R}^{\varepsilon}f^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}_{R},$
$\Gamma_{\bm{\mathcal{D}},B}^{(01),\ast}f=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\leavevmode\nobreak\ \delta_{1}\in\mathcal{E}^{1}\\ \delta_{1}\neq\varepsilon_{1}\end{subarray}}\frac{1}{\sqrt{|R_{1}|}}\left(h_{R_{1}}^{\delta_{1}}\otimes\frac{\mathbf{1}_{R_{2}}}{|R_{2}|}\right)B_{R}^{\varepsilon}f^{1\oplus\varepsilon_{1}\oplus\delta_{1},\varepsilon_{2}}_{R},$
therefore $(\Gamma_{\bm{\mathcal{D}},B}^{(10)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(10),\ast}$ as well as $(\Gamma_{\bm{\mathcal{D}},B}^{(01)})^{\ast}=\Gamma_{\bm{\mathcal{D}},B^{\ast}}^{(01),\ast}$ in the unweighted $\mathrm{L}^{2}(\mathbb{R}^{n+m};\mathbb{C}^{d})$ sense.
Two-matrix weighted bounds for the above mixed paraproducts can be easily deduced from two-matrix weighted bounds for the mixed type operators considered in [DKPS2024, Section 8]. Although at the time of [DKPS2024] only a rather incomplete treatment of the latter bounds was possible, they all follow now readily from Theorem 1 and Theorem 7. We give only one example below, the proof for the other operators being similar. Note that the parts of the proof that do not rely on the matrix weighted extension of the Fefferman–Stein vector valued inequalities were already carried out in [DKPS2024, Section 8]. Nevertheless, for the reader’s convenience we include full details.
@@
$[\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg{(}\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg{)}^{1/2},$
@@
To prove Lemma 12, we need a technical observation already present implicitly in the proof of [DKPS2024, Lemma 5.3].
@@
We use duality. Note that through an application of the monotone convergence theorem we may assume without loss of generality that only finitely many of the functions $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ are not identically equal to zero. Then, by the dyadic (unweighted) Littlewood–Payley estimates we have
@@
Let $g\in L^{p^{\prime}}(\mathbb{R}^{m})$ be arbitrary. Then, we have
@@
$\leq\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|g_{P}^{\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{L^{p^{\prime}}(\mathbb{R}^{m})}$
$\sim_{m,p}\bigg{\|}\bigg{(}\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg{)}^{1/2}\bigg{\|}_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\|g\|_{L^{p^{\prime}}(\mathbb{R}^{m})}.$
@@
Fix $x_{1}\in\mathbb{R}^{m}.$ For a.e. $x_{2}\in\mathbb{R}^{n},$ we denote by $\mathcal{W}_{x_{2},R_{1}}$ the reducing operator of the weight $W_{x_{2}}(y)\coloneq W(y,x_{2}),$ $y\in\mathbb{R}^{n}$ over any $R_{1}\in\mathcal{D}^{1}$ with respect to the exponent $p.$ For fixed $R_{1}\in\mathcal{D}^{1},$ we define $W_{R_{1}}(x_{2})\coloneq\mathcal{W}_{x_{2},R_{1}}^{p},$ for a.e. $x_{2}\in\mathbb{R},$ and denote by $\mathcal{W}_{R_{1},R_{2}}$ the reducing operator of $W_{R_{1}}$ over any $R_{2}\in\mathcal{D}^{2}$ with respect to the exponent $p.$ Applying now first (2.3), then part (1) of Lemma 3, and finally (2.4), we obtain
@@
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\frac{1}{p}}\bigg{\langle}\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg{\rangle}_{R_{2}},$
for all $\varepsilon_{2}\in\mathcal{E}^{2}$ and $R_{2}\in\mathcal{D}^{2}.$ Observe that in the last $\lesssim_{p,d},$ we used (2.4), as well as the fact that the supremum of the integrals is dominated by the integral of the supremum. So
@@
For a.e. $x_{2}\in\mathbb{R}^{m},$ using Theorem 7 in the first step and (2.6.3) in the second step, we obtain
@@
where we denote $W_{x_{1}}\coloneq W(x_{1},x_{2}),$ $x_{1}\in\mathbb{R}^{n}.$ Thus, applying Fubini–Tonelli again we get
@@
For a.e. $x_{1}\in\mathbb{R}^{n},$ applying first the matrix weighted bounds for the one-parameter dyadic square function from [Isralowitz_2020] and then (2.6.3), we get

### v3

@@ 4.2 Two-matrix weighted bounds for paraproducts
We now prove that $\|\Pi_{\bm{\mathcal{D}},B}^{(11)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ Using [14, Lemma 5.3] we have
@@
$b_{R}^{\varepsilon}\coloneq|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|,\quad R\in\bm{\mathcal{D}},\penalty 10000\ \varepsilon\in\mathcal{E}.$
By the well-known unweighted bounds for paraproducts in the scalar setting (see e. g. [3]) we have
@@
By definition, $\|b\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}}=\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$ Putting these facts together and the bound for the dyadic auxiliary biparameter maximal operator, we get
$\|\Pi_{\bm{\mathcal{D}},B}^{(11)}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m};\mathbb{C}^{d})}\lesssim\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|f\|_{\mathrm{L}^{p}(U)}.$
@@
$\|\Pi_{\bm{\mathcal{D}},B}^{(00)}\|_{\mathrm{L}^{p}(U)\rightarrow\mathrm{L}^{p}(V)}=\|\Pi_{\bm{\mathcal{D}},B^{\ast}}^{(11)}\|_{\mathrm{L}^{p^{\prime}}(V^{\prime})\rightarrow\mathrm{L}^{p^{\prime}}(U^{\prime})}\sim_{p,d}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}.$
We adapt the factorization trick from the proof of [23, Proposition 6.1]. We have
$(\Gamma_{\bm{\mathcal{D}},B}f,g)=\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}\langle B_{R}^{1\oplus\varepsilon\oplus\delta}f_{R}^{\varepsilon},g_{R}^{\delta}\rangle.$
@@
$|(\Gamma_{\bm{\mathcal{D}},B}f,g)|$ $=\bigg|\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}\frac{1}{\sqrt{|R|}}\text{tr}((B_{R}^{\varepsilon})^{\ast}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T})\bigg|$
@@
$\Phi_{R}^{\varepsilon}\coloneq\frac{1}{\sqrt{|R|}}\sum_{\begin{subarray}{c}\delta\in\mathcal{E}\\ \delta_{i}\neq\varepsilon_{i},\penalty 10000\ i=1,2\end{subarray}}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T},\quad R\in\bm{\mathcal{D}},\penalty 10000\ \varepsilon\in\mathcal{E}.$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\lesssim\|f\|_{\mathrm{L}^{p}(U)}\|g\|_{\mathrm{L}^{p^{\prime}}(V^{\prime})}.$
We have that $\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}$ is bounded by
$2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$\leq 2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{1\oplus\varepsilon\oplus\delta}|^{2}\cdot|\overline{f_{R}^{\delta}}^{T}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|^{2}}$
$=2^{n+m}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon,\delta\in\mathcal{E}\\ \varepsilon_{i}\neq\delta_{i},\penalty 10000\ i=1,2\end{subarray}}|V(x)^{-1/p}g_{R}^{\varepsilon}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\cdot|\mathcal{U}_{R}f_{R}^{\delta}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}$
@@
$\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}\leq 2^{n+m}\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m})}\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}.$
It is proved in [14, Lemma 5.2 and Corollary 5.4] that
$\|\widetilde{S}_{\bm{\mathcal{D}},U}f\|_{\mathrm{L}^{p}(\mathbb{R}^{n+m})}\lesssim\|f\|_{\mathrm{L}^{p}(U)},\qquad\|S_{\bm{\mathcal{D}},V^{\prime}}g\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{n+m})}\lesssim\|g\|_{\mathrm{L}^{p^{\prime}}(V^{\prime})},$
@@
Two-matrix weighted bounds for “mixed” paraproducts can be easily deduced from two-matrix weighted bounds for the mixed type operators considered in [14, Section 8]. Although at the time of [14] only a rather incomplete treatment of the latter bounds was possible, they all follow now readily from Theorem 1 and Theorem 6. To show how to apply our results, we give only one example below, the proof for the other operators being similar. Note that the parts of the proof that do not rely on the matrix weighted extension of the Fefferman–Stein vector valued inequalities were already carried out in [14, Section 8]. Nevertheless, for the reader’s convenience we include full details.
As our example, we focus on the mixed paraproduct
$\Gamma_{\bm{\mathcal{D}},B}^{(10)}f\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E},\penalty 10000\ \delta_{2}\in\mathcal{E}^{2}\\ \delta_{2}\neq\varepsilon_{2}\end{subarray}}\frac{1}{\sqrt{|R_{2}|}}h_{R}^{\varepsilon_{1},1\oplus\varepsilon_{2}\oplus\delta_{2}}B_{R}^{\varepsilon}\langle f^{\delta_{2},2}_{R_{2}}\rangle_{R_{1}},$
acting on suitable $\mathbb{C}^{d}$-valued functions $f$ on $\mathbb{R}^{n+m}.$ The notation and definitions for the other possible paraproducts can be found in the work of Holmes–Petermichl–Wick [23, Subsection 6.1], together with their duality relations.
@@
$\widetilde{M}\widetilde{S}]_{\bm{\mathcal{D}},U}f(x)\coloneq\bigg(\sum_{\begin{subarray}{c}R_{2}\in\mathcal{D}^{2}\\ \varepsilon_{2}\in\mathcal{E}^{2}\end{subarray}}(\sup_{R_{1}\in\mathcal{D}^{1}}|\mathcal{W}_{R}\langle f^{\varepsilon_{2},2}_{R_{2}}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1}))^{2}\frac{\mathbf{1}_{R_{2}}(x_{2})}{|R_{2}|}\bigg)^{1/2},$
@@
To prove Lemma 11, we need a technical observation already present implicitly in the proof of [14, Lemma 5.3].
@@
We use duality. As usual, by the monotone convergence theorem we may consider only finitely many of the functions $\{f_{P,\varepsilon}\}_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}$ to be nonzero. Then, by the dyadic (unweighted) Littlewood–Payley estimates we have
@@
Let $g\in\mathrm{L}^{p^{\prime}}(\mathbb{R}^{m})$ be arbitrary. Then, we have
@@
$\leq\bigg\|\bigg(\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg)^{1/2}\bigg\|_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\bigg\|\bigg(\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|g_{P}^{\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg)^{1/2}\bigg\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{m})}$
$\sim_{m,p}\bigg\|\bigg(\sum_{\begin{subarray}{c}P\in\mathcal{D}^{2}\\ \varepsilon\in\mathcal{E}^{2}\end{subarray}}|f_{P,\varepsilon}|^{2}\frac{\mathbf{1}_{P}}{|P|}\bigg)^{1/2}\bigg\|_{\mathrm{L}^{p}(\mathbb{R}^{m})}\cdot\|g\|_{\mathrm{L}^{p^{\prime}}(\mathbb{R}^{m})}.$
@@
Fix $x_{1}\in\mathbb{R}^{m}.$ For a.e. $x_{2}\in\mathbb{R}^{n},$ we denote by $\mathcal{W}_{x_{2},R_{1}}$ the reducing operator of the weight $W_{x_{2}}(y)\coloneq W(y,x_{2}),$ $y\in\mathbb{R}^{n}$ over any $R_{1}\in\mathcal{D}^{1}$ with respect to the exponent $p.$ For fixed $R_{1}\in\mathcal{D}^{1},$ we define $W_{R_{1}}(x_{2})\coloneq\mathcal{W}_{x_{2},R_{1}}^{p},$ for a.e. $x_{2}\in\mathbb{R},$ and denote by $\mathcal{W}_{R_{1},R_{2}}$ the reducing operator of $W_{R_{1}}$ over any $R_{2}\in\mathcal{D}^{2}$ with respect to the exponent $p.$ Let us observe here in general that if $\mathcal{V}_{E}$ is the reducing operator of a matrix weight $V$ over a set $E$ with respect to exponent $p,$ and $V^{-1/(p-1)}$ is also integrable over $E$ with reducing exponent $\mathcal{V}^{\prime}_{E}$ with respect to exponent $p^{\prime}$, then
$|\mathcal{V}_{E}e|\lesssim_{p,d}|\mathcal{V}_{E}\mathcal{V}_{E}^{\prime}|\cdot{\langle|V^{1/p}e|\rangle}_{E},\quad\forall e\in\mathbb{C}^{d};$ (4.16)
for the proof of this fact see, for instance, [14, Lemmas 3.1 and 3.2]. Applying now first (2.2), then (4.16), and finally (2.5), we obtain
@@
$\lesssim_{p,d}[W]_{A_{p},\bm{\mathcal{D}}}^{\frac{1}{p}}\bigg\langle\sup_{R_{1}\in\mathcal{D}^{1}}|W_{R_{1}}^{1/p}\langle f_{R_{2}}^{\varepsilon_{2},2}\rangle_{R_{1}}|\mathbf{1}_{R_{1}}(x_{1})\bigg\rangle_{R_{2}}$
for all $\varepsilon_{2}\in\mathcal{E}^{2}$ and $R_{2}\in\mathcal{D}^{2}.$ So we get
@@
For a.e. $x_{2}\in\mathbb{R}^{m},$ using Theorem 6 in the first step and (2.4) in the second step, we obtain
@@
where we denote $W_{x_{1}}\coloneq W(x_{1},x_{2}),$ $x_{1}\in\mathbb{R}^{n}.$ Fubini–Tonelli now yields
@@
For a.e. $x_{1}\in\mathbb{R}^{n},$ applying first the matrix weighted bounds for the one-parameter dyadic square function from [29] and then (2.4), we get